Bruno deposits $550 in an account that pays 8.3% simple interest.
How long will it take for Bruno's investment to double if he never
withdraws money from the account?
Ⓐ 13 months
Ⓑ 22 months
Ⓒ 145 months
Ⓓ 642 months

Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

simple interest formula is i = p * r * t
p = 550
r = .083 per year.
at simple interest, this becomes .083 per year / 12 = .0069166667 per month.

if you want to double your money, then i must be equal to 550.
formula becomes 550 = 550 * .0069166667 * t.
solve for t to get t = 550 / (550 * ,0069166667) = 144.5783133 months.

your closest selection to that is 145 months.
that's your solution.




Answer by greenestamps(13200)



Since the problem is multiple choice with only 4 choices (which are all very different from each other), the correct answer can be found by estimating.

The 8.3% interest is an annual interest rate.
8.3% is very close to 1/12; 12 times 8.3% is99.6%.
So the money will double (i.e., grow by 100%) in about 12 years, which is 144 months.

Only answer choice C is close to 144 months.

ANSWER: 145 months